OpenStudy (anonymous):

f(x)=x^2+6x-27 a. vertex. b. axis of symmetry. c. intercepts. d. graph. e. interval f is increasing. f. interval f is decreasing.

OpenStudy (jdoe0001):

\(\bf \textit{vertex of a parabola}\\ \quad \\ f(x) = {\color{red}{ 1}}x^2{\color{blue}{ +6}}x{\color{green}{ -27}} \quad \left(-\cfrac{{\color{blue}{ b}}}{2{\color{red}{ a}}}\quad ,\quad {\color{green}{ c}}-\cfrac{{\color{blue}{ b}}^2}{4{\color{red}{ a}}}\right)\) axis of symmetry, well, use the coordinates from the vertex intercepts, set f(x) = 0, that is \(\bf 0=x^2+6x-27\) solve for "x" like you'd any quadratic, by factoring graph, well, use the vertex, pick a point to the left of it, and one to the right of it and graph away where is it increasing or decreasing? well, look at the graph :)

OpenStudy (anonymous):

vertex: (-3,25.5) ? @jdoe0001

OpenStudy (jdoe0001):

well.... not quite

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

please explain?

OpenStudy (jdoe0001):

\(\bf \textit{vertex of a parabola}\\ \quad \\ f(x) = {\color{red}{ 1}}x^2{\color{blue}{ +6}}x{\color{green}{ -27}} \quad \left(-\cfrac{{\color{blue}{ 6}}}{2{\color{red}{ (1)}}}\quad ,\quad {\color{green}{ -27}}-\cfrac{{\color{blue}{ 6}}^2}{4{\color{red}{ (1)}}}\right)\)

OpenStudy (anonymous):

ooooh i didnt see the 2 by the six

OpenStudy (anonymous):

(-3,-36) ?

OpenStudy (jdoe0001):

well... ahemm \(\bf \textit{vertex of a parabola}\\ \quad \\ f(x) = {\color{red}{ 1}}x^2{\color{blue}{ +6}}x{\color{green}{ -27}} \quad \left(-\cfrac{{\color{blue}{ 6}}}{2{\color{red}{ (1)}}}\quad ,\quad {\color{green}{ -27}}-\cfrac{{\color{blue}{ 6}}^2}{4{\color{red}{ (1)}}}\right)\\ \quad \\\implies \left(-3\quad ,\quad -27-\cfrac{36}{4}\right) \implies (-3\quad ,\quad -27-9)\)

OpenStudy (anonymous):

ya and then you subtract 9 from -27 dont you?

OpenStudy (jdoe0001):

ohh yea... yes you do ... ahemm right.. so -36 correct

OpenStudy (jdoe0001):

the parabola is "x" based, meaning is vertically opening the leading term coefficient is positive, so it's going upwards